OpenStudy (anonymous):

Rachel deposited $5,960.32 into a savings account with an interest rate of 4.2% compounded twice a year. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

use compound formula \[\text{final amount}= p(1+ r/n)^{nt}\]

OpenStudy (anonymous):

\[9000=5960.32(1+\frac{(.042)}{2})^{2*t}\]

OpenStudy (anonymous):

I just can't seem to understand how to find how many years/months it takes. I understand the formula though.

OpenStudy (mr.math):

You need to take ln of both sides and solve for t.

OpenStudy (mertsj):